OpenStudy (anonymous):

Can somebody please help/guide me through 2 of these questions? 1. For which pair of functions is (g of x)(a)= lal-2? A. f(a)=a^2-4 and g(a)= square root of a B. f(a)=1/2a-1 and g(a)=2a-2 C. f (a)=5+a^2 and g(a)= square root of a-5-2 D. f(a)=3-3a and g(a)=4a-5 2. Let f(x) be a polynomial function of degree 2 and g(x) be a polynomial function of degree 3. What is the degree of (f of g)(x)? A.2 B.3 C.5 D.6 I got 6 is that correct?

OpenStudy (phi):

For the 2 question, do you mean What is the degree of (f o g)(x)? (not f of g)(x) f o g means replace x in f with g(x) g(x) is polynomial of degree 3. example x^3 + x^2 + x +1 f(x) is degree 2. example x^2 + x + 1 when we put g(x) in for x in f(x) we get this mess (x^3 + x^2 + x +1)^2 + (x^3 + x^2 + x +1) + 1 but luckily all we care about is the highest exponent, which will come from x^3 * x^3 = x^6 so yes, 6 is the answer.

OpenStudy (anonymous):

thank you! and yes that's what I meant.

OpenStudy (phi):

did you try Q1?

OpenStudy (anonymous):

Nope :(

OpenStudy (phi):

(g o f)(a) means replace all "a" s in g with f(a) test the choices. write down g(a): g= sqr(a) replace "a" with f(a): sqr(a^2-4) a^2 -4 is not a perfect square... it factors into (a-2)(a+2) so that is not the answer can you try the second choice?

OpenStudy (anonymous):

It's C.

OpenStudy (phi):

yes, I think so, but it depends on what g(a)= square root of a-5-2 means if it means \[ g(a) = \sqrt{a-5} - 2 \] then when we replace a with f(a) = 5+a^2 we get \[ g(a) = \sqrt{5+a^2-5} - 2 \\= \sqrt{a^2} - 2 = a-2\] because we take the positive root, we should write it as |a| - 2

OpenStudy (anonymous):

Thanks! :)

OpenStudy (phi):

yw